solve: 1/y^2+5/y+3=0
Answer by Alan3354(69443)

solve: 1/y^2+5/y+3=0
-----------
Multiply by y^2
1 + 5y + 3y^2 = 0
3y^2 + 5y + 1 = 0
Can you finish it?